OpenStudy (anonymous):

y1 -3/y^2 = 3/x^2 would be y1 = -y^2/x^2 ?

OpenStudy (anonymous):

can't wait to see this answer ...

OpenStudy (freckles):

\[y' \cdot \frac{-3}{y^2}=\frac{3}{x^2} ?\] and you are solving for y'?

OpenStudy (freckles):

multiply both sides by the reciprocal of what is next to the y'

OpenStudy (freckles):

and yes

OpenStudy (freckles):

He was doing derivative junk earlier so I think he meant y' instead of y1

OpenStudy (anonymous):

yes

OpenStudy (freckles):

and I think he was doing the question find y' given 1/x+1/y=1

OpenStudy (anonymous):

ok thanks just making sure